OpenStudy (anonymous):

x divided by 3 over 6/3x

OpenStudy (debbieg):

\[\Large \dfrac{ \dfrac{ x }{ 3 } }{ \dfrac{ 6 }{ 3x } }\]

OpenStudy (debbieg):

divide by a fraction, by multiplying by the reciprocal. Change it to multiplication.....